INFRARED SPECTROSCOPY AND ULTRA-VIOLET PHOTOCHEMISTRY OF H_{2}O_{2} IN SOLID PARAHY-DROGEN

\underline{RICHARD L. COLE}, SHAWN M. PALMER, and DAVID T. ANDERSON, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

The high resolution infrared spectrum of solid parahydrogen crystals doped with anhydrous H_{2}O_{2} has been observed for the first time. Doped crystals were prepared via the rapid vapor deposition method developed by Fajardo and co-workers. UHP (urea hydrogen peroxide) was utilized as a source of anhydrous H_{2}O_{2} vapor. A series of photolysis studies were conducted with 274 nm radiation generated from the AS_{2} (second anti-Stokes) beam of a 355 nm pumped hydrogen Raman shifter. Mechanistic studies will be presented as well as infrared spectroscopic studies of the photochemically generated radical species.